Below is data for two variables, X and Y. Answer the questions that follow.

X Y X-MX Y-MY (X-MX)2 (Y-MY)2 (X-MX)(Y-MY)
0 4 -3 -1 9 1 3
2 3 -1 -2 1 4 ___
4 7 1 ___ 1 4 2
6 6 3 1 ___ 1 3

MX = _____; MY = _____

MX = 5; MY = 4

MX = 3; MY = 5

MX = 4; MY = 4

MX = 12; MY = 20

What is the missing value under the Y-MY column?

4

-2

2

3

What is the missing value under the (X-MX)2 column?

3

-3

5

9

What is the missing value under the (X-MX)(Y-MY) column? (far right column)

-3

2

-2

3

SSX = _____; SSY = _____

SSX = 10; SSY = 10

SSX = 10; SSY = 20

SSX = 100; SSY = 400

SSX = 20; SSY = 10

SP = ____

20

100

12

10

r = ____

-.71

+.85

-.85

+.71

Positive or negative correlation?

Negative

Positive
Answer #1

Foto Givene X Y X - MX Y-MYCX-Mx)?|CY - My)? 74-6x) (۱۲۰ (۱۶) 4 9 3. 2 3 - 1 - 2 1 2 4 4 - 1 2 2 4 6 7 6 3 9 3. - 12 20 20 10(x-mx (3) g (x. - Mx Cr. My - 2 = (-1) (-2) SSX E (X-MX) ssy & cr- My)² xss = 20 SSY OTE REDMI NOTE PRI MI DUAL CAMERA ds E CRO Yz Ell X - My) (Y- My) Icessx) (554) tro (20) y 0.71Result :- حكم (1 3 ; My blo y Y - My 2 3) (x-mx ² = 9 4) CX-MX2 CY- My) 3 5) ssx = 20 ID SsY = 6> Spela 7) 0.71 positive corr